OpenStudy (anonymous):

A truck accelerates from rest to a speed of 17 m/s in 9.1 s. If the tires have a radius of 33 cm, how many revolutions do the tires make during this time

OpenStudy (anonymous):

http://www.physics247.com/physics-homework-help/rotational-motion.php this should help

OpenStudy (shamim):

i wanna give u some formula for solvinh this math

OpenStudy (shamim):

a=(v-u)/t v^2=u^2+2*a*s N=s/(2*pi*r)

OpenStudy (isaiah.feynman):